[A] Consider the inner product space obtained by equipping ?[0,2] with the inner product given below:
〈?(?),?(?)〉 = ∫ ?(?)?(?)?? 2 0


Determine the value of each of the following (simplifying where possible; no decimals). You do not have to show the steps of calculating the integrals, but must at least write the integrals used in your calculations.   

(A.1) 〈?,1〉   




(A.2) ‖ ? − 1 ‖

(A.3) ?(??,?? + 10), i.e. the distance between ?? and ?? + 10 .   







(A.4) Determine the value of ? so that the functions ?(?) = ? and ?(?) = ?? + 2 are  orthogonal relative to the given inner product.

(A.5) Consider the subspace ℙ1 of ?[0,2] consisting of all polynomials of degree 1 or less. Use the Gram-Schmidt process to convert its standard basis ? = {1,?} to an orthonormal basis.

(A.6) Obtain the polynomial ?(?) ∈ ℙ1 that “best approximates” ?(?) = ?2 in the sense that ?(?(?),?(?)) is minimized. Express your final answer in the form ?(?) = ?? + ? where ? and ? are simplified as much as possible (no decimal approximations).
Answer #1

A.1) Using the definition of inner product given we have

\left \langle t,1 \right \rangle=\int_{0}^{2} (t)(1)dt =\left ( t^2\over 2 \right )_{0}^{2}=2

A.2) As  \| x\| =\sqrt\left \langle x,x \right \rangle so \|t-1\|= \sqrt\left \langle t-1,t-1 \right \rangle

Now,

2 t - 1)2 t - 1,t-1) = (t - - 1) (t - 1) dt = یا | نت 3

Therefore, \|t-1\|= \sqrt {2\over 3}

A-3) We know d(x,y)=\| y-x\|

So,

d(e^t,e^{t+10})=\| e^{t+10} - e^{t} \| \\ ~~~~~~~~~~~~~~~~~~~= |e^{10}-1| \| e^{t} \| \\ ~~~~~~~~~~~~~~~~~~~=|e^{10}-1| \sqrt{\left \langle e^{t}, e^{t} \right \rangle}

Now,

{\left \langle e^{t}, e^{t} \right \rangle}=\int_{0}^{2}(e^{t})(e^t)dt \\ ~~~~~~~~~~~~~~=\left ( e^{2t}\over 2 \right )_{0}^{2}\\ ~~~~~~~~~~~~~~={1\over 2}[e^{4}-1]

Therefore, d(e^{t}, e^{t+10})=|e^{10}-1|\sqrt{{1\over 2}[e^{4}-1]}

A-4) f(t) and g(t) are orthogonal if \left\langle f(t),g(t) \right \rangle=0

Using given definition we get

\left\langle f(t),g(t) \right \rangle= \int_{0}^{2}(t)(at+2) dt \\ ~~~~~~~~~~~~~~~~~~~=\left ( a{t^3\over 3}+t^2 \right )_{0}^{2}\\ ~~~~~~~~~~~~~~~~~~~={8\over 3}a+4

Thus, \left\langle f(t),g(t) \right \rangle=0 gives

{8\over 3}a+4=0 \\ \Rightarrow a=-{3\over 2}

NOTE: As per rules first 4 parts are solved at a time